Out of 8000 graduates in a town 800 are females, out of 1600

graduate employees 120 are females. Use 2  at 5% level to

determine if any distinction is made in appointment on the basis

of sex:


1
Expert's answer
2021-01-18T09:02:51-0500
"\begin{matrix}
 \text{Number of} & \text{Number of} & \text{Total number of} \\
 \text{Female\ \ \ \ \ \ } & \text{male\ \ \ \ \ \ \ \ \ \ } & \text{employees\ \ \ \ \ \ \ \ \ } \\
\text{employees} & \text{employees} & \text{} 
\end{matrix}""\begin{matrix}
 800 \ \ \ \ \ \ \ \ \ \ \ & 7200\ \ \ \ \ \ \ \ \ \ \ & 8000 \ \ \ \ \ \ \ \ \ \ \ \\
 120 \ \ \ \ \ \ \ \ \ \ \ & 1480\ \ \ \ \ \ \ \ \ \ \ & 1600 \ \ \ \ \ \ \ \ \ \ \ \\
 920 \ \ \ \ \ \ \ \ \ \ \ & 8680\ \ \ \ \ \ \ \ \ \ \ & 9600 \ \ \ \ \ \ \ \ \ \ \ \\
\end{matrix}"


"\dfrac{920\times8000}{9600}=766.7"




"\dfrac{8680\times8000}{9600}=7233.3"

"\dfrac{920\times1600}{9600}=153.3"

"\dfrac{8680\times1600}{9600}=1446.7"

"\begin{matrix}
 \text{Number of} & \text{Number of} & \text{Total number of} \\
 \text{Female\ \ \ \ \ \ } & \text{male\ \ \ \ \ \ \ \ \ \ } & \text{employees\ \ \ \ \ \ \ \ \ } \\
\text{employees} & \text{employees} & \text{} 
\end{matrix}""\begin{matrix}
 766.7 \ \ \ \ \ \ \ \ \ \ & 7233.3 \ \ \ \ \ \ \ \ \ \ \ & 8000 \ \ \ \ \ \ \ \ \ \ \ \ \ \ \\
 153.3 \ \ \ \ \ \ \ \ \ \ & 1446.7\ \ \ \ \ \ \ \ \ \ \ & 1600 \ \ \ \ \ \ \ \ \ \ \ \ \ \ \\
 920 \ \ \ \ \ \ \ \ \ \ \ \ & 8680\ \ \ \ \ \ \ \ \ \ \ \ \ \ & 9600 \ \ \ \ \ \ \ \ \ \ \ \ \ \ \\
\end{matrix}"

Observed value (O)

Expected value(E)


"\def\arraystretch{1.5}
 \begin{array}{c:c:c:c :c}
 O & E & O-E & (O-E)^2 & \dfrac{(O-E)^2}{E} \\ \hline
 800 & 766.7 & 33.3 & 1108.9 & 1.45 \\
120 & 153.3 & -33.3 & 1108.9 & 7.23 \\
7200 & 7233.3 & 33.3 & 1108.9 & 0.15 \\
1480 & 1446.7 & 33.3 & 1108.9 & 0.77 \\
 \hdashline
 & & & Total & 9.60
\end{array}"

Degree of freedom "=(R-1)(C-1)=(2-1)(2-1)=1"

The critical value of "\chi^2" at "\alpha=0.05" for "1" d.f. "=3.841"

Since the calculated value of "\chi^2" "(9.60)" is greater than the critical value of "\chi^2" at "\alpha=0.05" for "1" d.f. "(3.841)," then the null hypothesis is rejected and the alternative hypothesis is accepted.

Therefore there is a distinction is made in appointment on the basis of sex.